You just took over another financial adviser's account. The client invested $15,500 at 6.92% compounded monthly and now has $24,980.58. How long has this client had the money invested?
Answer #1

Future Value = Present Value(1+r)t

Let the no of period be x

24980.58 = $ 15500(1+0.0692)x

(1+0.0692)x = 24980.58 /15500

(1+0.0692)x = 1.61165

Takinbg log both sides

log (1+0.0692)x = log 1.61165

x log 1.0692 = log 1.61165

taking value of log

x * 0.0291 = 0.2073

x = 0.2072/ 0.0291

x = 7.12 Months

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.